OpenStudy (anonymous):

(X^2-7x+12)/(x^2-2x-3) Can someone simplify this for me? Any help would definitely be appreciated.

OpenStudy (tester97):

http://prntscr.com/3923xg use this

OpenStudy (ankitshaw):

factorising x^2-7x+12 = (x-3)(x-4) and x^2-2x-3 = (x-3)(x+1) so (x-3) gets cut we have (x-3)/(x+1) as answer

OpenStudy (anonymous):

I appreciate your help so much, thank you.

OpenStudy (ankitshaw):

sorry its (x-4)/(x+1)